OpenStudy (anonymous):

A horse race has 11 entries. Assuming that there are no ties, what is the probability that the two horses owned by one person finish first and second?

OpenStudy (campbell_st):

P(win) = 2/11 then the owner has 1 horse to run second from the remaining 10 horses P(2nd) = 1/10 P(1st and 2nd) = 2/11 x 1/10